OpenStudy (anonymous):

what are the steps to reduce (64sin^2(pi/6))/64

OpenStudy (campbell_st):

start by eliminating the denominator  next find the exact value of \[\sin(\frac{\pi}{6}) \]  then its simply using the exact value \[(\sin(\frac{\pi}{6}))^2\]